# Exploring specialization and sensitivity of convolutional neural networks in the context of simultaneous image augmentations

### v1

@@ 3.2.2 Single-class sensitivity to augmentations
Tables 2, 3 list the top-5 classes sensitive to considered augmentations according to conducted sensitivity analysis. Corresponding selectivity patterns are problematic to be deduced by divergent values only. Thus, we matched sensible classes and top-5 masked predictions by means of the Jaccard index (Tables S7.6(a)–(f)). Table 4 is provided separately since it contains the most prominent evidence of proper sensible class selection manifested as a substantial increase of Jaccard indices for erasing augmentation applied to AlexNet’s input. The objects of corresponding classes (screen; monitor; television; see Table 3) contain a filled rectangular pattern, similar to the results of applying the erasing transform. Other notable examples are classes containing light-shadow or edge contrast claimed as sensitive to contrast adjustment (cliff dwelling; jigsaw puzzle). Detailed analysis of other classes and their augmentation sensitivity is provided in D.
@@
Erasing (.1) | rock python (62/.58); magnetic compass (635/.54); paperknife (623/.54); maraca (641/.52); ladle (618/.51); | Hungarian pointer (211/.89); centipede (79/.88); poke bonnet (452/.84); thatched roof (853/.69); Doberman (236/.68); | beaker (438/.49); hussar monkey (371/.46); swimming cap (433/.45); ski mask (796/.44); drilling platform (540/.44); |
Sharp. () | church (497/.29); banister (421/.28); bee eater (92/.28); leaf beetle (304/.28); overskirt (689/.27); | ram, tup (348/.48); Pomeranian (259/.41); vestment (887/.38); kimono (614/.38); triumphal arch (873/.36); | poncho (735/.32); harvester, reaper (595/.28); Windsor tie (906/.27); face powder (551/.26); grasshopper (311/.25); |
Rolling () | mashed potato (935/.50); platypus (103/.48); cardigan (474/.47); cradle (516/.45); head cabbage (936/.42); | Scotch terrier (199/.67); Australian terrier (193/.66); harmonica (593/.65); chain saw (491/.61); barracouta (389/.59); | chocolate sauce (960/.50); English springer spaniel (217/.40); neck brace (678/.38); rock crab (119/.38); trilobite (69/.37); |
Grayscale () | Brittany spaniel (215/.34); tennis ball (852/.34); police van (734/.34); sea slug (115/.33); Welsh springer spaniel (218/.33); | Holocanthus tricolor (392/.56); American lobster (122/.53); briard (226/.52); African crocodile (49/.46); salamandra (25/.46); | coral reef (973/.39); wig (903/.32); Irish setter (213/.32); lakeside (975/.30); napkin (529/.30); |
Gaus.blur () | mask (643/.50); dingo (273/.46); European gallinule (136/.42); coho salmon (391/.40); soup bowl (809/.38); | pencil box (709/.61); eel (390/.57); harvester (595/.52); stole (824/.52); panda (388/.51); | affenpinscher (252/.44); vestment (887/.44); broom (462/.43); green snake (55/.38); hair slide (584/.36); |
Bright. () | hatchet (596/.37); Indian cobra (63/.35); miniskirt (655/.35); Gordon setter (214/.35); plate rack (729/.33); | howler monkey, howler (379/.53); plate (923/.49); worm fence (912/.49); boxer (242/.46); china cabinet (495/.44); | bookshop (454/.49); stove (827/.40); cliff, drop (972/.36); scabbard (777/.32); king snake (56/.31); |
Contrast. () | napkin (529/.45); pig (341/.39); basenji (253/.37); malinois (225/.34); African crocodile (49/.33); | vestment (887/.51); parking meter (704/.49); jigsaw puzzle (611/.48); gyromitra (993/.48); waffle iron (891/.46); | Irish terrier (184/.41); Australian terrier (193/.36); drumstick (542/.35); wing (908/.34); holster (597/.34); |
Satur. () | holster (597/.40); refrigerator (760/.37); boa constrictor (61/.35); whiptail lizard (41/.33); rotisserie (766/.31); | screen (782/.65); wombat (106/.59); rifle (764/.49); water bottle (898/.46); three-toed sloth (364/.44); | nipple (680/.38); dough (961/.38); sandbar (977/.37); walking stick (313/.36); electric switch (844/.34); |
Hue () | bluetick (164/.41); spaghetti squash (940/.39); hen-of-the-woods (996/.38); bloodhound (163/.37); boxer (242/.37); | wok (909/.54); Lakeland terrier (189/.46); fox squirrel (335/.46); jeep (609/.45); orangutan (365/.44); | chiffonier (493/.62); pizza (963/.38); radiator (753/.36); soft-coated wheaten terrier (202/.35); paddle wheel (694/.35); |
Hflip () | conch (112/.32); plunger (731/.26); hatchet (596/.26); oxcart (690/.25); limousine (627/.24); | lab coat (617/.39); cheetah (293/.38); bull mastiff (243/.37); drum (541/.35); cocktail shaker (503/.32); | carbonara (959/.25); blowfish (397/.23); groom (982/.22); horizontal bar (602/.22); Norfolk terrier (185/.22); |
Rotation () | solar collector (807/.48); basset hound (161/.37); punch bag (747/.36); sliding door (799/.36); breakwater (460/.36); | platypus (103/.59); buckle (464/.54); face powder (551/.51); paddlewheel (694/.50); bathtub (435/.49); | hotpot (926/.36); shoji (789/.33); hyena (276/.31); weasel (356/.30); abacus (398/.30); |
El.loc.blur () | hippopotamus (344/.52); albatross (146/.48); Sussex spaniel (220/.48); water snake (58/.46); paperknife (623/.46); | scabbard (777/.76); dishwasher (534/.65); computer mouse (673/.65); wreck (913/.58); mosque (668/.57); | cheeseburger (933/.55); dough (961/.52); Chesapeake Bay retriever (209/.50); whiskey jug (901/.50); toy terrier (158/.47); |
@@
Erasing (.1) | screen (782/.06); television (851/.05); monitor (664/.04); lighter (626/.03); pedestal (708/.03); | warthog (343/.17); tree frog (31/.10); scuba diver (983/.07); safe (771/.05); fiddler crab (120/.03); | wardrobe (894/.04); frilled lizard (43/.03); television (851/.02); monitor (664/.02); screen (782/.02); |
Sharp. () | affenpinscher (252/.03); rubber (767/.02); drumstick (542/.01); steel drum (822/.01); Great Dane (246/.01); | warthog (343/.11); coho salmon (391/.10); ladle (618/.03); Alaskan malamute (249/.02); carton (478/.02); | wardrobe (894/.05); overskirt (689/.02); lighthouse (437/.02); jackfruit (955/.02); gazelle hound (176/.01); |
Rolling () | Appenzeller (240/.02); vault (884/.01); cabbage butterfly (324/.01); grand piano (579/.01); drilling platform (540/.01); | black grouse (80/.04); space shuttle (812/.04); ricksha (612/.03); grocery store (582/.03); cliff dwelling (500/.02); | English springer spaniel (217/.03); flat-coated retriever (205/.02); toilet seat (861/.01); strawberry (949/.01); German short-haired pointer (210/.01); |
Grayscale () | tree frog (31/.04); affenpinscher (252/.03); Shetland sheepdog (230/.03); missile (657/.03); television (851/.03); | warthog (343/.16); tree frog (31/.10); scuba diver (983/.07); space shuttle (812/.05); Russian wolfhound (169/.04); | tree frog (31/.02); sweatshirt (841/.02); lifeboat (625/.02); chain (488/.02); Shetland sheepdog (230/.02); |
Gaus.blur () | wastebin (412/.03); affenpinscher (252/.03); tub, vat (876/.03); bucket (463/.02); beaker (438/.02); | flat-coated retriever (205/.08); scuba diver (983/.07); stove (827/.06); bubble (971/.06); coho salmon (391/.06); | three-toed sloth (364/.03); English springer spaniel (217/.03); hare (331/.02); sleeping bag (797/.02); lighthouse (437/.02); |
Bright. () | Weimaraner (178/.04); mailbox (637/.04); koala (105/.03); frilled lizard (43/.02); pipe organ (687/.02); | Bedlington terrier (181/.08); pinwheel (723/.03); Weimaraner (178/.03); grey wolf (269/.03); barrel, cask (427/.03); | upright piano (881/.02); ambulance (407/.02); cellular phone (487/.01); Psittacus erithacus (87/.01); hippopotamus (344/.01); |
Contrast. () | walking stick (313/.03); red-breasted merganser (98/.03); pay-phone (707/.02); chain (488/.02); sock (806/.02); | cliff dwelling (500/.12); jigsaw puzzle (611/.11); cornet (513/.06); shovel (792/.05); snow leopard (289/.04); | jigsaw puzzle (611/.05); convertible (511/.03); bubble (971/.02); spotlight, spot (818/.01); hook, claw (600/.01); |
Satur. () | hay (958/.06); cliff dwelling (500/.03); indris (384/.02); moped (665/.02); sea urchin (328/.02); | cliff dwelling (500/.21); wardrobe (894/.08); sea urchin (328/.04); barbell (422/.04); grey wolf (269/.04); | jellyfish (107/.02); gown (578/.02); bubble (971/.02); upright piano (881/.02); sturgeon (394/.01); |
Hue () | Weimaraner (178/.04); agaric (992/.03); missile (657/.03); Chihuahua (151/.03); vine snake (59/.02); | cliff dwelling (500/.20); house finch (12/.18); salamandra (25/.07); breastplate (461/.06); missile (657/.04); | bee (309/.03); cabbage butterfly (324/.02); china cabinet (495/.02); digital clock (530/.02); jellyfish (107/.01); |
Hflip () | giant schnauzer (197/.02); missile (657/.02); shield (787/.02); moped (665/.02); snowmobile (802/.02); | cliff dwelling (500/.16); shield (787/.07); Siberian husky (250/.04); snowmobile (802/.03); missile (657/.03); | Granny Smith (948/.03); nipple (680/.02); fur coat (568/.02); combination lock (507/.02); sturgeon (394/.01); |
Rotation () | jeep (609/.03); napkin (529/.03); red-breasted merganser (98/.03); Dandie Dinmont terrier (194/.02); leafhopper (317/.02); | husky (248/.07); ruffed grouse (82/.06); bullfrog (30/.06); beaker (438/.04); Siberian husky (250/.04); | tray (868/.04); cellular phone (487/.03); military uniform (652/.03); mushroom (947/.02); coyote (272/.01); |
El.loc.blur () | sport car (817/.02); golf cart (575/.02); bulletproof vest (465/.02); black and gold garden spider (72/.02); cup (968/.01); | cliff dwelling (500/.20); barrel, cask (427/.07); parking meter (704/.07); coyote (272/.03); cornet (513/.03); | apron (411/.03); cleaver (499/.02); lacewing fly (318/.01); stingray (6/.01); odometer (685/.01); |

### v2

@@ 3.2.2 Single-class sensitivity to augmentations
Tables 2–4 list the top-5 classes sensitive to considered augmentations according to conducted sensitivity analysis. Corresponding selectivity patterns are problematic to be deduced by divergent values only. Thus, we matched sensitive classes and top-5 masked predictions by means of the Jaccard index (Tables S7.6.1 (a)–(f), S7.6.2 (a)–(f)) and highlighted the outliers according to the precomputed threshold values (see D). Table 5 is provided separately since it contains the most prominent evidence of proper sensitive class selection manifested as a substantial increase of Jaccard indices for erasing augmentation applied to AlexNet’s input. The objects of corresponding classes (screen; monitor; television; see Table 3) contain a filled rectangular pattern, similar to the results of applying the erasing transform. Other notable examples are classes containing light-shadow or edge contrast claimed as sensitive to contrast adjustment (cliff dwelling; jigsaw puzzle). Detailed analysis of other classes and their augmentation sensitivity is provided in E.
@@
| Erasing | rock python (62/.58) | Hungarian pointer (211/.89) | beaker (438/.49) |
| (.1) | magnetic compass (635/.54) | centipede (79/.88) | hussar monkey (371/.46) |
| paperknife (623/.54) | poke bonnet (452/.84) | swimming cap (433/.45) | |
| maraca (641/.52) | thatched roof (853/.69) | ski mask (796/.44) | |
| ladle (618/.51) | Doberman (236/.68) | drilling platform (540/.44) | |
| Sharp. | church (497/.29) | ram, tup (348/.48) | poncho (735/.32) |
| () | banister (421/.28) | Pomeranian (259/.41) | harvester, reaper (595/.28) |
| bee eater (92/.28) | vestment (887/.38) | Windsor tie (906/.27) | |
| leaf beetle (304/.28) | kimono (614/.38) | face powder (551/.26) | |
| overskirt (689/.27) | triumphal arch (873/.36) | grasshopper (311/.25) | |
| Rolling | mashed potato (935/.50) | Scotch terrier (199/.67) | chocolate sauce (960/.50) |
| () | platypus (103/.48) | Australian terrier (193/.66) | English spaniel (217/.40) |
| cardigan (474/.47) | harmonica (593/.65) | neck brace (678/.38) | |
| cradle (516/.45) | chain saw (491/.61) | rock crab (119/.38) | |
| head cabbage (936/.42) | barracouta (389/.59) | trilobite (69/.37) | |
| Grayscale | Brittany spaniel (215/.34) | Holocanthus tricolor (392/.56) | coral reef (973/.39) |
| () | tennis ball (852/.34) | American lobster (122/.53) | wig (903/.32) |
| police van (734/.34) | briard (226/.52) | Irish setter (213/.32) | |
| sea slug (115/.33) | African crocodile (49/.46) | lakeside (975/.30) | |
| Welsh spaniel (218/.33) | salamandra (25/.46) | napkin (529/.30) | |
| Gaus.blur | mask (643/.50) | pencil box (709/.61) | affenpinscher (252/.44) |
| () | dingo (273/.46) | eel (390/.57) | vestment (887/.44) |
| European gallinule (136/.42) | harvester (595/.52) | broom (462/.43) | |
| coho salmon (391/.40) | stole (824/.52) | green snake (55/.38) | |
| soup bowl (809/.38) | panda (388/.51) | hair slide (584/.36) | |
| Bright. | hatchet (596/.37) | howler monkey, howler (379/.53) | bookshop (454/.49) |
| () | Indian cobra (63/.35) | plate (923/.49) | stove (827/.40) |
| miniskirt (655/.35) | worm fence (912/.49) | cliff, drop (972/.36) | |
| Gordon setter (214/.35) | boxer (242/.46) | scabbard (777/.32) | |
| plate rack (729/.33) | china cabinet (495/.44) | king snake (56/.31) | |
| Contrast. | napkin (529/.45) | vestment (887/.51) | Irish terrier (184/.41) |
| () | pig (341/.39) | parking meter (704/.49) | Australian terrier (193/.36) |
| basenji (253/.37) | jigsaw puzzle (611/.48) | drumstick (542/.35) | |
| malinois (225/.34) | gyromitra (993/.48) | wing (908/.34) | |
| African crocodile (49/.33) | waffle iron (891/.46) | holster (597/.34) | |
| Satur. | holster (597/.40) | screen (782/.65) | nipple (680/.38) |
| () | refrigerator (760/.37) | wombat (106/.59) | dough (961/.38) |
| boa constrictor (61/.35) | rifle (764/.49) | sandbar (977/.37) | |
| whiptail lizard (41/.33) | water bottle (898/.46) | walking stick (313/.36) | |
| rotisserie (766/.31) | three-toed sloth (364/.44) | electric switch (844/.34) | |
| Hue | bluetick (164/.41) | wok (909/.54) | chiffonier (493/.62) |
| () | spaghetti squash (940/.39) | Lakeland terrier (189/.46) | pizza (963/.38) |
| hen-of-the-woods (996/.38) | fox squirrel (335/.46) | radiator (753/.36) | |
| bloodhound (163/.37) | jeep (609/.45) | wheaten terrier (202/.35) | |
| boxer (242/.37) | orangutan (365/.44) | paddle wheel (694/.35) | |
| Hflip | conch (112/.32) | lab coat (617/.39) | carbonara (959/.25) |
| () | plunger (731/.26) | cheetah (293/.38) | blowfish (397/.23) |
| hatchet (596/.26) | bull mastiff (243/.37) | groom (982/.22) | |
| oxcart (690/.25) | drum (541/.35) | horizontal bar (602/.22) | |
| limousine (627/.24) | cocktail shaker (503/.32) | Norfolk terrier (185/.22) | |
| Rotation | solar collector (807/.48) | platypus (103/.59) | hotpot (926/.36) |
| () | basset hound (161/.37) | buckle (464/.54) | shoji (789/.33) |
| punch bag (747/.36) | face powder (551/.51) | hyena (276/.31) | |
| sliding door (799/.36) | paddlewheel (694/.50) | weasel (356/.30) | |
| breakwater (460/.36) | bathtub (435/.49) | abacus (398/.30) | |
| El.loc.blur | hippopotamus (344/.52) | scabbard (777/.76) | cheeseburger (933/.55) |
| () | albatross (146/.48) | dishwasher (534/.65) | dough (961/.52) |
| Sussex spaniel (220/.48) | computer mouse (673/.65) | Ches.Bay retriever (209/.50) | |
| water snake (58/.46) | wreck (913/.58) | whiskey jug (901/.50) | |
| paperknife (623/.46) | mosque (668/.57) | toy terrier (158/.47) |
@@
| Erasing | screen (782/.06) | warthog (343/.17) | wardrobe (894/.04) |
| (.1) | television (851/.05) | tree frog (31/.10) | frilled lizard (43/.03) |
| monitor (664/.04) | scuba diver (983/.07) | television (851/.02) | |
| lighter (626/.03) | safe (771/.05) | monitor (664/.02) | |
| pedestal (708/.03) | fiddler crab (120/.03) | screen (782/.02) | |
| Sharp. | affenpinscher (252/.03) | warthog (343/.11) | wardrobe (894/.05) |
| () | rubber (767/.02) | coho salmon (391/.10) | overskirt (689/.02) |
| drumstick (542/.01) | ladle (618/.03) | lighthouse (437/.02) | |
| steel drum (822/.01) | Alaskan malamute (249/.02) | jackfruit (955/.02) | |
| Great Dane (246/.01) | carton (478/.02) | gazelle hound (176/.01) | |
| Rolling | Appenzeller (240/.02) | black grouse (80/.04) | English spaniel (217/.03) |
| () | vault (884/.01) | space shuttle (812/.04) | flat-coated retriever (205/.02) |
| cabbage butterfly (324/.01) | ricksha (612/.03) | toilet seat (861/.01) | |
| grand piano (579/.01) | grocery store (582/.03) | strawberry (949/.01) | |
| drilling platform (540/.01) | cliff dwelling (500/.02) | German short-haired pointer (210/.01) | |
| Grayscale | tree frog (31/.04) | warthog (343/.16) | tree frog (31/.02) |
| () | affenpinscher (252/.03) | tree frog (31/.10) | sweatshirt (841/.02) |
| Shetland sheepdog (230/.03) | scuba diver (983/.07) | lifeboat (625/.02) | |
| missile (657/.03) | space shuttle (812/.05) | chain (488/.02) | |
| television (851/.03) | Russian wolfhound (169/.04) | Shetland sheepdog (230/.02) | |
| Gaus.blur | wastebin (412/.03) | flat-coated retriever (205/.08) | three-toed sloth (364/.03) |
| () | affenpinscher (252/.03) | scuba diver (983/.07) | English spaniel (217/.03) |
| tub, vat (876/.03) | stove (827/.06) | hare (331/.02) | |
| bucket (463/.02) | bubble (971/.06) | sleeping bag (797/.02) | |
| beaker (438/.02) | coho salmon (391/.06) | lighthouse (437/.02) | |
| Bright. | Weimaraner (178/.04) | Bedlington terrier (181/.08) | upright piano (881/.02) |
| () | mailbox (637/.04) | pinwheel (723/.03) | ambulance (407/.02) |
| koala (105/.03) | Weimaraner (178/.03) | cellular phone (487/.01) | |
| frilled lizard (43/.02) | grey wolf (269/.03) | Psittacus erithacus (87/.01) | |
| pipe organ (687/.02) | barrel, cask (427/.03) | hippopotamus (344/.01) | |
| Contrast. | walking stick (313/.03) | cliff dwelling (500/.12) | jigsaw puzzle (611/.05) |
| () | red-breasted merganser (98/.03) | jigsaw puzzle (611/.11) | convertible (511/.03) |
| pay-phone (707/.02) | cornet (513/.06) | bubble (971/.02) | |
| chain (488/.02) | shovel (792/.05) | spotlight, spot (818/.01) | |
| sock (806/.02) | snow leopard (289/.04) | hook, claw (600/.01) | |
| Satur. | hay (958/.06) | cliff dwelling (500/.21) | jellyfish (107/.02) |
| () | cliff dwelling (500/.03) | wardrobe (894/.08) | gown (578/.02) |
| indris (384/.02) | sea urchin (328/.04) | bubble (971/.02) | |
| moped (665/.02) | barbell (422/.04) | upright piano (881/.02) | |
| sea urchin (328/.02) | grey wolf (269/.04) | sturgeon (394/.01) | |
| Hue | Weimaraner (178/.04) | cliff dwelling (500/.20) | bee (309/.03) |
| () | agaric (992/.03) | house finch (12/.18) | cabbage butterfly (324/.02) |
| missile (657/.03) | salamandra (25/.07) | china cabinet (495/.02) | |
| Chihuahua (151/.03) | breastplate (461/.06) | digital clock (530/.02) | |
| vine snake (59/.02) | missile (657/.04) | jellyfish (107/.01) | |
| Hflip | giant schnauzer (197/.02) | cliff dwelling (500/.16) | Granny Smith (948/.03) |
| () | missile (657/.02) | shield (787/.07) | nipple (680/.02) |
| shield (787/.02) | Siberian husky (250/.04) | fur coat (568/.02) | |
| moped (665/.02) | snowmobile (802/.03) | combination lock (507/.02) | |
| snowmobile (802/.02) | missile (657/.03) | sturgeon (394/.01) | |
| Rotation | jeep (609/.03) | husky (248/.07) | tray (868/.04) |
| () | napkin (529/.03) | ruffed grouse (82/.06) | cellular phone (487/.03) |
| red-breasted merganser (98/.03) | bullfrog (30/.06) | military uniform (652/.03) | |
| Dandie Dinmont terrier (194/.02) | beaker (438/.04) | mushroom (947/.02) | |
| leafhopper (317/.02) | Siberian husky (250/.04) | coyote (272/.01) | |
| El.loc.blur | sport car (817/.02) | cliff dwelling (500/.20) | apron (411/.03) |
| () | golf cart (575/.02) | barrel, cask (427/.07) | cleaver (499/.02) |
| bulletproof vest (465/.02) | parking meter (704/.07) | lacewing fly (318/.01) | |
| black and gold garden spider (72/.02) | coyote (272/.03) | stingray (6/.01) | |
| cup (968/.01) | cornet (513/.03) | odometer (685/.01) |
| Shapley values | Sobol indices | |||
| AlexNet | ResNet18 | AlexNet | ResNet18 | |
| Erasing | orchard (249/.53) | ice cream parlor (185/.36) | home theater (177/.07) | home theater (177/.08) |
| (.1) | temple asia (330/.36) | lagoon (204/.34) | movie theater indoor (235/.04) | television room (328/.05) |
| village (348/.33) | archaelogical excavation (13/.30) | television room (328/.04) | movie theater indoor (235/.03) | |
| greenhouse indoor (165/.33) | gymnasium indoor (168/.29) | phone booth (263/.03) | server room (298/.02) | |
| beer garden (53/.31) | corral (105/.29) | barndoor (41/.03) | barndoor (41/.02) | |
| Sharp. | delicatessen (114/.18) | delicatessen (114/.14) | motel (231/.02) | harbor (171/.02) |
| () | bazaar outdoor (47/.16) | veterinarians office (346/.14) | field wild (141/.01) | church outdoor (91/.01) |
| excavation (136/.15) | garage indoor (156/.13) | building facade (67/.01) | gazebo exterior (159/.01) | |
| art school (20/.14) | islet (194/.13) | ball pit (34/.01) | elevator lobby (130/.01) | |
| lawn (209/.14) | tundra (341/.13) | canal urban (79/.01) | temple asia (330/.01) | |
| Rolling | field road (142/.42) | underwater ocean deep (342/.29) | throne room (331/.02) | church outdoor (91/.02) |
| () | canyon (81/.38) | hotel outdoor (181/.28) | galley (155/.01) | ballroom (35/.01) |
| rice paddy (287/.36) | tree house (339/.25) | nursery (240/.01) | rope bridge (291/.01) | |
| dining room (121/.26) | greenhouse outdoor (166/.24) | bar (39/.01) | pond (271/.01) | |
| wave (357/.25) | natural history museum (239/.24) | movie theater indoor (235/.01) | aquarium (9/.01) | |
| Grayscale | mountain path (233/.33) | topiary garden (333/.25) | hospital room (179/.04) | park (254/.02) |
| () | lawn (209/.30) | general store indoor (160/.23) | archive (14/.03) | church outdoor (91/.02) |
| sandbox (294/.26) | swimming pool outdoor (326/.23) | music studio (238/.03) | swimming pool outdoor (326/.02) | |
| flea market indoor (146/.25) | crevasse (111/.22) | coffee shop (99/.03) | pet shop (261/.01) | |
| florist shop indoor (147/.25) | florist shop indoor (147/.21) | physics laboratory (264/.03) | playroom (269/.01) | |
| Gaus.blur | market indoor (222/.31) | courtyard (109/.29) | throne room (331/.02) | coffee shop (99/.02) |
| () | childs room (89/.24) | playroom (269/.26) | shed (299/.02) | barndoor (41/.02) |
| shoe shop (300/.23) | basement (43/.25) | barndoor (41/.02) | server room (298/.01) | |
| driveway (127/.23) | fabric store (137/.22) | temple asia (330/.02) | shed (299/.01) | |
| kitchen (203/.22) | patio (259/.22) | restaurant (284/.02) | ballroom (35/.01) | |
| Bright. | pizzeria (267/.30) | dressing room (126/.26) | childs room (89/.03) | rock arch (289/.02) |
| () | patio (259/.22) | corral (105/.21) | playroom (269/.03) | bazaar indoor (46/.02) |
| general store outdoor (161/.21) | pier (266/.20) | creek (110/.02) | berth (55/.02) | |
| crosswalk (112/.21) | locker room (219/.20) | candy store (80/.02) | playroom (269/.02) | |
| childs room (89/.19) | loading dock (216/.20) | art studio (21/.02) | waiting room (352/.02) | |
| Contrast. | village (348/.26) | archaelogical excavation (13/.28) | television studio (329/.02) | desert vegetation (117/.02) |
| () | balcony exterior (32/.24) | army base (18/.28) | music studio (238/.02) | television studio (329/.02) |
| promenade (273/.23) | utility room (343/.24) | childs room (89/.02) | kindergarden classroom (202/.01) | |
| biology laboratory (56/.20) | veterinarians office (346/.23) | amusement arcade (6/.01) | pasture (258/.01) | |
| creek (110/.20) | volleyball court outdoor (351/.23) | stage indoor (315/.01) | pet shop (261/.01) | |
| Satur. | archaelogical excavation (13/.26) | garage outdoor (157/.28) | discotheque (122/.06) | ice shelf (187/.04) |
| () | balcony interior (33/.25) | kindergarden classroom (202/.23) | ice shelf (187/.04) | discotheque (122/.03) |
| restaurant kitchen (285/.23) | beer hall (54/.22) | stage indoor (315/.04) | stage indoor (315/.03) | |
| trench (340/.23) | bookstore (60/.22) | ball pit (34/.04) | candy store (80/.03) | |
| wet bar (358/.21) | sauna (295/.20) | ice floe (186/.03) | ice floe (186/.03) | |
| Hue | recreation room (281/.25) | kasbah (200/.37) | botanical garden (62/.03) | vegetable garden (345/.03) |
| () | slum (308/.24) | landfill (206/.30) | vegetable garden (345/.03) | florist shop indoor (147/.03) |
| ruin (292/.24) | rainforest (279/.27) | stage indoor (315/.03) | beer hall (54/.02) | |
| beer hall (54/.24) | fishpond (145/.26) | florist shop indoor (147/.03) | discotheque (122/.02) | |
| plaza (270/.23) | fastfood restaurant (139/.25) | rainforest (279/.02) | botanical garden (62/.02) | |
| Hflip | sandbox (294/.15) | residential neighborhood (283/.17) | basement (43/.01) | japanese garden (197/.01) |
| () | general store indoor (160/.14) | zen garden (364/.14) | forest road (152/.01) | hospital room (179/.01) |
| fishpond (145/.13) | catacomb (85/.13) | candy store (80/.01) | cottage (107/.01) | |
| trench (340/.12) | diner outdoor (119/.11) | church indoor (90/.01) | market outdoor (223/.01) | |
| locker room (219/.12) | food court (148/.11) | staircase (317/.01) | aqueduct (10/.01) | |
| Rotation | driveway (127/.28) | office (244/.22) | street (319/.02) | office (244/.01) |
| () | boat deck (58/.27) | village (348/.22) | golf course (164/.01) | cottage (107/.01) |
| campus (77/.26) | oast house (242/.20) | ticket booth (332/.01) | vegetable garden (345/.01) | |
| clothing store (96/.23) | ski slope (305/.20) | atrium public (25/.01) | football field (149/.01) | |
| bakery shop (31/.23) | home office (176/.19) | field cultivated (140/.01) | florist shop indoor (147/.01) | |
| El.loc.blur | galley (155/.43) | hangar outdoor (170/.33) | beer hall (54/.03) | lecture room (210/.02) |
| () | nursing home (241/.37) | desert sand (116/.33) | forest road (152/.01) | youth hostel (363/.02) |
| entrance hall (134/.32) | desert vegetation (117/.31) | rice paddy (287/.01) | hospital room (179/.02) | |
| kitchen (203/.29) | elevator lobby (130/.31) | stage outdoor (316/.01) | dining hall (120/.01) | |
| general store indoor (160/.29) | residential neighborhood (283/.28) | rope bridge (291/.01) | greenhouse indoor (165/.01) | |
@@
| Augmentation of input | AlexNet | ResNet18 | ||||||||
| q=0.5 | q=0.6 | q=0.7 | q=0.8 | q=0.9 | q=0.5 | q=0.6 | q=0.7 | q=0.8 | q=0.9 | |
| Erasing | .0888 | .1101 | .1412 | .2 | .2434 | .0249 | .0316 | .0334 | .0434 | .0672 |
| Sharp. | .005 | .0052 | .0054 | .0056 | .0056 | .0096 | .0093 | .0095 | .0092 | .0089 |
| Rolling | .0075 | .0079 | .0083 | .008 | .0067 | .0065 | .0091 | .0081 | .0054 | .0042 |
| Grayscale | .0232 | .022 | .0209 | .0202 | .0191 | .003 | .0028 | .0032 | .0042 | .0041 |
| Gaus.blur | .0076 | .008 | .0088 | .0096 | .0091 | .0119 | .0143 | .0135 | .0085 | .0074 |
| Bright. | .0239 | .0265 | .0267 | .0297 | .0218 | .0004 | .0004 | .0007 | .0012 | .0034 |
| Contrast. | .037 | .0473 | .0639 | .0907 | .0845 | .0006 | .0007 | .0008 | .001 | .0021 |
| Satur. | .0246 | .029 | .0332 | .0598 | .0726 | .0452 | .0638 | .0764 | .0748 | .0617 |
| Hue | .0287 | .033 | .0382 | .0579 | .0563 | .057 | .0749 | .0829 | .0785 | .0652 |
| Hflip | .0051 | .0055 | .0057 | .0059 | .0075 | .0017 | .0038 | .0085 | .0089 | .0091 |
| Rotation | .0064 | .0065 | .0068 | .0085 | .0092 | .0048 | .0055 | .0059 | .0063 | .0065 |
| El.loc.blur | .0051 | .0055 | .0056 | .006 | .0059 | .0023 | .0044 | .0045 | .0039 | .004 |